4. Which combination of acid and base would you use to make a buffer system that has a pH of 7.72 (5pts) Show sample work.


Acid Conjugate Base H SO H SO: HOHO, HF HCIO CIO C. HOH CHO Ka 1.5x102 2.4x10-12 3.5x10-4 3.0x10- 1.3x10-10
Answer #1

4) Heo and co- Ka = 3:0x10-8 pH = -lof 3.0 - log 10-8 = -0.4771 +8 TpH = 7.522) which is near about 7.7 5 (6th OH & GGO- Ka =by using [PH = pka = -logka ), we will calculate the pH of the buffers. 1) And 14.503 4503 115x10 pH = pka = -lopka ConjugateA buffer solution is a mixture is a mixture of a weak acid and its conjugate base or a weak base and its conjugate acid. It helps to maintain a near constant pH upon the addition of small amounts of H+ or OH- ions to a solution.

HClO and ClO- combination of acid and base would be used to make a buffer system that has a pH of near about 7.7.

It can be calculated by using Henderson-Hasselbalch equation which is eplained in the following images: